OpenStudy (anonymous):

Complete the square to find the vertex. y^2+6y+8x+1=0 I have completed the square and got to this point: (y+3)^2=8x-1+9 I know the "y" cordinate of the vertex is -3 but how do I find the "x"?

OpenStudy (anonymous):

some mistake here i think

OpenStudy (anonymous):

\[(y+3)^2=-8x-1+9\]or \[(y+3)^2=-8x+8\]

OpenStudy (anonymous):

rewrite as \[(y+3)^2=-8(x-1)\] and eyeball the vertex as \((1,-3)\)

OpenStudy (anonymous):

Now I see it! I forgot to combine like terms. Thank you!

OpenStudy (anonymous):

yw you also need \(-8x\) on the right hand side, not \(8x\)